Problem 2.7 A loan of $8,000 must be repaid with 6 year-end level payments (i.e., constant pay- ments). The effective annual loan rate is 11%. What is the annual payment? Problem 2.8 You make a deposit now into an account earning 6% annually in return for a payment of 250 at the end of each of the next 8 years. What should you deposit today? Problem 2.9 An annuity immediate has semi-annual payments of 1,000 for 25 years at a rate of 6%, convertible quarterly. Find the present value.
Answer #1

2.7 PV = 8000
Loan Rate = 11%
Number of Periods(n) = 6
Annual Payment PMT = 8000/(1-(1+11%)-6)/11% = 1,891.01

2.8 PMT = 250
Number of Periods = 8
Rate = 6%
PV = 250*(1-(1+6%)-8)/6% = 1552.45